Bob’s lawn mowing service is a profit maximizing firm operating in a perfectly competitive

market. Bob mows lawns for $27 each. His total cost each day is $280, or which $30 is a fixed cost. He mows 10 lawns a day. What are Bob’s decision rules about when to shut down and when to exit the market? Answer using numbers! Refer to the shut down and exit rules from the lectures on perfect competition.










1
Expert's answer
2022-02-17T14:44:56-0500

When to Shut-down and when to Exit

"TC=280"

"FC=30"

"VC=TC-FC"

"VC=280-30=250"

"TR_{daily}=\$27 \times 10\ lawns=\$270"

"Profit(\pi)=TR-TC"

"\pi (day\ one)=270-280=-10"

From above, we see that Bob's short-run revenue is able to finance variable costs. Bob is in a position to continue business for the time covered by the fixed cost. Assuming the fixed cost contract expires in a month;


"TR_{month}=(\$27 \times 10\ lawns)\times30=\$8100"

"VC_{month}=(\$250\times 30)=\$7500"

"FC_{month}=\$ 30"

"TC_{month}=7500+30=\$7530"

"\pi\ per\ month=8100-7530=\bold{\$570}"

In the long run, Bob is able to finance business operations expenses and earn a profit.